what is the height of a triangle that has an area of 60 yard squared and a base with a length of 12 yards
Answer by TutorDelphia(193)

When you have a problem involving a formula, first write it down
A=(1/2)b*h
Next plug in the known values into the formula
60=(1/2)12*h
60=6*h
divide both sides by 6
10=h